OpenStudy (tester97):

Divide –2x³ + 21x² – 20x – 12 by (x – 3).  (2 points)

OpenStudy (anonymous):

x^3__x^2__x^1__x^0 | 3 ---------------------------------------... -2___21__-20__-12 ____-6___45___75 -2__15___25___63 Therefore: -2x^3 + 21x^2 -20x -12 = (x-3) * ( -2x^2 + 15x + 25) + 63 2. Let y = (3 - 4x)^2 , this function's domain is the set of real numbers and its range is the set of all non-negative numbers. f-1 exists if for any given y in f's range, solving for x gives a unique x: y = (3 - 4x)^2 <=> sqrt(y) = abs(3 - 4x) <=> <=> 3 - 4x = sqrt(y) OR 3 - 4x = -sqrt(y) <=> <=> x = [ 3 - sqrt(y) ] / 4 OR x = [ 3 + sqrt(y) ] / 4 Since for every y>0 x is not well defined, f-1 does not exist (x is well defined only for y = 0)

OpenStudy (tester97):

what?